# Cleaning data with Swipe

### v1

@@ Abstract
The repair problem for functional dependencies is the problem where an input database needs to be modified such that all functional dependencies are satisfied and the difference with the original database is minimal. The output database is then called an optimal repair. If the allowed modifications are value updates, finding an optimal repair is $\mathsf{NP}$-hard. A well-known approach to find approximations of optimal repairs builds a Chase tree in which each internal node resolves violations of one functional dependency and leaf nodes represent repairs. A key property of this approach is that controlling the branching factor of the Chase tree allows to control the trade-off between repair quality and computational efficiency. In this paper, we explore an extreme variant of this idea in which the Chase tree has only one path. To construct this path, we first create a partition of attributes such that classes can be repaired sequentially. We repair each class only once and do so by fixing the order in which dependencies are repaired. This principle is called priority repairing and we provide a simple heuristic to determine priority. The algorithms for attribute partitioning and priority repair are combined in the Swipe algorithm. An empirical study on four real-life data sets shows that Swipe is in the range of one to three orders of magnitude faster than multi-sequence Chase-based approaches, whereas the quality of repairs is comparable or better. Moreover, we provide a scalability analysis of the Swipe algorithm.
@@ 1 Introduction
Although we can easily find a minimal-cost repair in Example 1, it is known that the repair problem for FDs as defined here, cannot be solved efficiently. Deciding whether there exists a repair with a cost lower than a constant $C$, is $\mathsf{NP}$-complete and finding constant-factor approximations of minimal-cost repairs is $\mathsf{NP}$-hard [30]. To deal with this, algorithms have been proposed to search for approximate solutions [8, 30, 5, 32]. In the current paper, we focus on one particular approach that relies on the Chase algorithm and has been implemented [^1] We note that Llunatic can in fact deal with Equality Generating Dependencies (EGDs), which subsume FDs. in the open-source framework Llunatic [24, 25]. Originally, the Chase algorithm was designed to work as a proof engine for dependency implication [3, 1]. Informally, the algorithm takes as input (i) a set of tuples and (ii) a set of FDs and then builds a sequence of Chase steps. In each such Chase step, one FD that is currently violated on the given set of tuples, becomes satisfied by equating the right-hand side attributes of the FD. The output of a Chase step is then used as input for the next Chase step and this process continues until all FDs are satisfied. In the original algorithm, tuples contained labelled variables and conflict resolution among variables with different labels is done by a consistent choice (e.g., if labels are integer-valued indices, we can always choose the variable with the lower index). In this setting, it is well-known that the Chase algorithm terminates and satisfies the Church Rosser property [1]. The latter property states that the output of the Chase algorithm does not depend on the order in which FDs are selected.
@@
Because it is infeasible to generate the complete Chase tree, Llunatic uses a cost manager that allows the definition of a wide range of pruning strategies. Examples of these strategies include (i) limiting the number of outgoing branches of a node, (ii) limiting the amount of leaf nodes and (iii) limiting modifications to forward repairs only [25]. The cost manager thus offers a trade-off between repair quality (i.e., generating more repairs implies a higher chance of a finding a high quality repair) and computational cost (i.e., generating more repairs implies larger Chase trees).
@@ 5.1 Data sets
The study involves four real-life data sets, which differ in size, number of errors and number of considered constraints. Two of them (Hospital and Flight) are commonly used in experiments on data quality, while the others have been composed and used in previous work by the authors. Each data set is accompanied by a gold standard that contains the correct tuples for a sample of the data set.
@@
  • Hospital. A benchmark data set often used in literature on data cleaning [16, 36]. The original data [16] stems from the U.S. Department of Health and Human Services [^6] http://www.hospitalcompare.hhs.gov. We use here the truncated version [36]. The running example from Figure 1 is a sample from this data set.
@@ 5.2 Repair techniques
We compare the Swipe algorithm with Llunatic [25]. The data sets were stored in a PostgreSQL database in separate schemas. Note that we used PostgreSQL version 9.5 because Llunatic relies on the ‘‘with oids’’ option during table creation, which is abandoned as of version 12999See https://www.postgresql.org/docs/12/release-12.html, section E.13.2.
@@
Llunatic is a general purpose Chase engine that builds a Chase tree where nodes correspond to (sequences of) repair steps and leaf nodes represent repairs. The implementation allows configuration of (i) pruning strategies to keep the search efficient and (ii) a partial order that indicates preferences of values for repair. We use the most recent version (2.0) available at GitHub [^10] https://github.com/donatellosantoro/Llunatic. We made a local build of the Llunatic code in order to be able to run it from the command line on an lxc container. We encoded the available FDs for each data set as EGDs in separate configuration files. As Llunatic allows for a wide range of configuration options, we decided to create two configurations for each data set. The first configuration (S) is a standard one and is the same for each data set. It uses the standard cost manager and searches for forward repairs only. We set both the branching threshold and the potential solutions threshold to $2$. This means that the Chase tree is binary and we stop searching after two repairs have been found. The partial order to resolve conflicts is based on the frequency of values. The main idea of this standard configuration is that it uses a general strategy to resolve conflicts (majority voting) and has quite strong pruning capabilities. The second configuration is a fine-tuned configuration (FT) that is tailored to each data set individually as it uses a more refined partial order and a more relaxed pruning strategy. The details of this second configuration are given for each data set in Table 3.
Swipe. The Swipe algorithm was implemented in Java (Version 8 or higher) as part of the $\mathsf{ledc}$ framework [^11] https://gitlab.com/ledc/ledc-fundy, which is an open-source framework for data quality. The implementation contains a set of five basic repair functions: majority vote ($\mathsf{mv}$), choice of the maximal value ($\mathsf{max}$), choice of the minimal value ($\mathsf{min}$) and two weighted voting functions ($\mathsf{wv_{1}}$ and $\mathsf{wv_{2}}$). With the first voting function $\mathsf{wv_{1}}$, weights are assigned to values in the value bag by a weight encoded in a dedicated weighting attribute. Such a function is for example useful if we can determine a weight based on the age of a value, source of the value… The second voting function $\mathsf{wv_{2}}$ generates a weight based on the number of null values that occur in the row where a value is taken. This voting system, where the weight of a value in the value bag depends on the row in which that value is found, has been used in several other repair systems [30, 10]. For each data set, we chose a suitable repair function and used it for all attributes in that data set (see Table 3).
Table 3: Configuration details of LLunatic (FT) and Swipe for individual data sets.
Data set Swipe ($\rho_{a}$) Llunatic (FT)
Hospital $\mathsf{mv}$
@@
Allergen $\mathsf{max}$ Preference for higher values
EudraCT $\mathsf{wv_{2}}$
@@ 5.3 Precision and recall
Table 4: Overview of precision ($P$), recall ($R$) and $F$-score obtained for the different repair methods. Best results are marked in bold font.
Data set Llunatic (S) Llunatic (FT) Swipe
$P$ Hospital [0.93,0.96] [0.11,0.12] 0.96
Allergen [0.00,0.57] [0.64,0.64] 0.64
Eudract [0.80,0.84] [0.80,0.84] 0.84
Flight [0.67,0.72] [0.80,0.80] 0.77
$R$ Hospital [0.83,0.85] [0.29,0.31] 0.89
Allergen [0.00,0.54] [0.30,0.30] 0.30
Eudract [0.31,0.32] [0.31,0.32] 0.37
Flight [0.66,0.71] [0.54,0.54] 0.75
$F$ Hospital [0.88,0.90] [0.16,0.17] 0.92
Allergen [0.00,0.55] [0.41,0.41] 0.41
Eudract [0.44,0.47] [0.44,0.47] 0.52
Flight [0.67,0.72] [0.65,0.65] 0.76
Table 4 provides an overview of the repair quality obtained by the two configurations of Llunatic and Swipe on all data sets. For each combination of data set and measure of quality, we indicate the best performing approach in bold font. For recall and $F$-score on the Allergen data set and for precision on the Hospital data set, we did not indicate a best approach as it cannot be indicated with certainty. When inspecting the results in Table 4, several interesting observations can be made.
A first observation is that the results of Llunatic are sensitive to the exact configurations. For example, for Allergen, we encoded a partial order that enforces the same behavior of the $\mathsf{max}$ repair function and this configuration provides much better results than the standard approach where the partial order is based on frequency. In fact, it was confirmed that in this case, Llunatic and Swipe produced exactly the same results. For the Hospital data set, there is also a large difference in repair quality between the two configurations, most likely due to the usage of backward repairing in the second configuration. In general, we found that giving Llunatic the possibility to include backward repairs (this is the case in the fine-tuned configurations) seems to decrease repair quality on the data sets considered here. Generalizing this observation should be done with caution, but the results reported here make a strong case for at least questioning the usefulness of backward repair on real-life data sets.
@@ 5.4 Scalability
In this section, we investigate the run time efficiency and scalability of Swipe. First, we measure the mean run time over five runs of Swipe and Llunatic on the different data sets and report the mean execution times (in seconds). We do not report other scalability parameters like CPU usage or main memory consumption, although we have investigated these parameters and came to similar conclusions. In case of Llunatic, we differentiate between the standard configuration and the configuration that was fine-tuned for each data set (Table 3). The results are reported in Table 5.
@@
Data set Llunatic (S) Llunatic (FT) Swipe
@@
The results from Table 5 show that Swipe outperforms Llunatic on each data set in terms of execution time. This comes as no surprise as Swipe is designed explicitly as a simplification of the multi-sequence Chase algorithm that underlies Llunatic. Yet, we observe that the gain in execution time is considerable. Swipe is two to three orders of magnitude faster than the fine-tuned configuration of Llunatic. When we compare to the standard Llunatic configuration, the difference reduces to one order of magnitude. This provides us with some interesting insights. First of all, the cost manager of Llunatic does what it is supposed to do: restricting repair operations to forward repair only and limiting the branching factor significantly reduces the execution time of Llunatic significantly. However, Table 4 shows that this gain in computational efficiency also leads to lesser quality of the repairs. In that sense, the Swipe algorithm makes more considerate choices in reducing computational cost with better repair quality as a result.

### v2

@@ Abstract
The repair problem for functional dependencies is the problem where an input database needs to be modified such that all functional dependencies are satisfied and the difference with the original database is minimal. The output database is then called an optimal repair. If the allowed modifications are value updates, finding an optimal repair is $\mathsf{NP}$-hard. A well-known approach to find approximations of optimal repairs builds a Chase tree in which each internal node resolves violations of one functional dependency and leaf nodes represent repairs. A key property of this approach is that controlling the branching factor of the Chase tree allows to control the trade-off between repair quality and computational efficiency. In this paper, we explore an extreme variant of this idea in which the Chase tree has only one path. To construct this path, we first create a partition of attributes such that classes can be repaired sequentially. We repair each class only once and do so by fixing the order in which dependencies are repaired. This principle is called priority repairing and we provide a simple heuristic to determine priority. The techniques for attribute partitioning and priority repair are combined in the Swipe algorithm. An empirical study on four real-life data sets shows that Swipe is one to three orders of magnitude faster than multi-sequence Chase-based approaches, whereas the quality of repairs is comparable or better. Moreover, a scalability analysis of the Swipe algorithm shows that Swipe scales well in terms of an increasing number of tuples.
@@ 1 Introduction
Although we can easily find a minimal-cost repair in Example 1, it is known that the repair problem for FDs as defined here, cannot be solved efficiently. Deciding whether there exists a repair with a cost lower than a constant $C$, is $\mathsf{NP}$-complete and finding constant-factor approximations of minimal-cost repairs is $\mathsf{NP}$-hard [30]. To deal with this, algorithms have been proposed to search for approximate solutions [8, 30, 5, 32]. In the current paper, we focus on one particular approach that relies on the Chase algorithm and has been implemented in the open-source framework Llunatic [24, 25]. We note and emphasize here that Llunatic can in fact deal with constraints that are more expressive than FDs. Yet, we limit the explanation of Llunatic here to FDs only, as these constraints are the focus of the current paper. Originally, the Chase algorithm was designed to work as a proof engine for dependency implication [3, 1]. Informally, the algorithm takes as input (i) a set of tuples and (ii) a set of FDs and then builds a sequence of Chase steps. In each such Chase step, one FD that is currently violated on the given set of tuples, becomes satisfied by equating the right-hand side attributes of the FD. The output of a Chase step is then used as input for the next Chase step and this process continues until all FDs are satisfied. In the original algorithm, tuples contained labelled variables and conflict resolution among variables with different labels is done by a consistent choice (e.g., if labels are integer-valued indices, we can always choose the variable with the lower index). In this setting, it is well-known that the Chase algorithm terminates and satisfies the Church Rosser property [1]. The latter property states that the output of the Chase algorithm does not depend on the order in which FDs are selected.
@@
Because it is infeasible to generate the complete Chase tree, Llunatic uses a cost manager that allows the definition of a wide range of pruning strategies. Examples of these strategies include (i) limiting the number of outgoing branches of a node by means of a branching threshold, (ii) limiting the amount of leaf nodes by a potential solutions threshold and (iii) limiting modifications to forward repairs only [25]. The cost manager thus offers a trade-off between repair quality (i.e., generating more repairs implies a higher chance of a finding a high quality repair) and computational cost (i.e., generating more repairs implies larger Chase trees).
@@ 5.1 Data sets
The study involves four real-life data sets, which differ in size, number of errors and number of considered constraints. Two of them (Hospital [16, 36, 25] and Flight [31, 25, 10]) are commonly used in experiments on data quality, while the others have been composed and used in previous work by the authors. Each data set is accompanied by a gold standard that contains the correct tuples for a sample of the data set.
@@
  • Hospital. A benchmark data set often used in literature on data cleaning [16, 36]. The original data [16] stems from the U.S. Department of Health and Human Services [^5] http://www.hospitalcompare.hhs.gov. We use the version consisting of 1000 tuples that has been used before in data repairing experiments [36]. The running example from Figure 1 is a sample from this data set.
@@ 5.2 Repair techniques
We compare the Swipe algorithm with Llunatic [25]. The data sets were stored in a PostgreSQL database in separate schemas. Note that we used PostgreSQL version 9.5 because Llunatic relies on the ‘‘with oids’’ option during table creation, which is abandoned as of version 12888See https://www.postgresql.org/docs/12/release-12.html, section E.13.2.
Llunatic.
@@
Llunatic is a general purpose Chase engine that builds a Chase tree where nodes correspond to (sequences of) repair steps and leaf nodes represent repairs. The implementation allows configuration of (i) pruning strategies to keep the search efficient and (ii) a partial order that indicates preferences of values for repair. We use the most recent version (2.0) available at GitHub [^9] https://github.com/donatellosantoro/Llunatic. We made a local build of the Llunatic code in order to be able to run it from the command line on an lxc container. We encoded the available FDs for each data set in separate configuration files. As Llunatic allows for a wide range of configuration options, we decided to create two configurations for each data set:
  • The first configuration is the standard one (S) and is the same for each data set. It searches for forward repairs only and we set both the branching threshold and the potential solutions threshold to $2$. This means that the Chase tree is binary and we stop searching after two repairs have been found. The partial order to resolve conflicts is based on the frequency of values. The main idea of this standard configuration is that it resembles the search strategy of Swipe. More specifically, it resolves conflicts in the same way as the default method in Swipe (i.e., majority voting), it uses forward repairing only and it constructs few sequences. The main difference is thus that Swipe will construct a specific order on the FDs and uses no variables.
  • The second configuration is a fine-tuned configuration (FT) that is tailored to each data set individually. First of all, we allowed backward repairing for Hospital, Eudract and Flight to see if this helps in finding better repairs. We did not allow this for Allergen as there we wanted to focus on mirroring the behaviour of Swipe when using the $\mathsf{max}$ repair function. Second, to account for typographical errors that are present in the attributes of Hospital and Flight, we used a cost manager that uses Smith-Waterman string similarity between values to guide resolution. For Hospital, Eudract and Flight, we also increased the branching threshold such that a larger part of the search space is explored and more repairs are searched for.
A summary of the main differences of the FT configurations is given for each data set in Table 3. The full configuration files are available online and can be used for reproduction purposes 4.
Table 3: Configuration details of LLunatic (FT) for each data set.
Data set Llunatic (FT)
Hospital
@@
Allergen Mimic $\mathsf{max}$ repair function
Eudract
@@
Swipe.
The Swipe algorithm was implemented in Java (Version 8 or higher) as part of the $\mathsf{ledc}$ framework [^10] https://gitlab.com/ledc/ledc-fundy, which is an open-source framework that bundles various techniques for data quality under one umbrella project. The implementation of Swipe contains a set of basic repair functions. We consider three of them in our experiments:
  • The first one is a majority voting function ($\mathsf{mv}$) with random tie breaking. This repair function was used in the examples throughout this paper and is considered a default choice.
  • The second one is a weighted voting function ($\mathsf{wv}$) with random tie breaking. The weight of a value in the bag is based on the number of null values that occur in the row where that value is taken. More specifically, if a value from the bag is taken from a tuple with $N$ null values, the weight for that value is equal to $\left(|\mathcal{R}|-N\right)^{4}$. We include this repair function as it has been shown effective in previous work [10].
  • The third repair function simply chooses the maximal value ($\mathsf{max}$) and can be used in a scenario where attribute domains are equipped with an order relation that is relevant to the repair. This is true for all datasets included in the experiment, except for the Hospital dataset, where using the natural order simply sorts the attribute values alphabetically.
For each data set, we run Swipe three times and in each run use a different repair function. That repair function is then used for all attributes in the data set. Note that each of these three functions is preservative, which means Proposition 2 applies in each scenario we test here.
@@ 5.3 Precision and recall
Table 4: Overview of precision ($P$), recall ($R$) and $F$-score obtained for Llunatic (two configurations) and Swipe (three repair functions). Best results are marked in bold font. n/a indicates the repair function was not applicable on that dataset.
Data set Llunatic (S) Llunatic (FT) Swipe ($\mathsf{mv}$) Swipe ($\mathsf{wv}$) Swipe ($\mathsf{max}$)
$P$ Hospital [0.93,0.96] [0.11,0.12] 0.96 0.96 n/a
Allergen [0.00,0.57] [0.64,0.64] 0.21 0.21 0.64
Eudract [0.80,0.84] [0.80,0.84] 0.83 0.84 0.23
Flight [0.67,0.72] [0.80,0.80] 0.71 0.79 0.56
$R$ Hospital [0.83,0.85] [0.29,0.31] 0.89 0.89 n/a
Allergen [0.00,0.54] [0.30,0.30] 0.10 0.10 0.30
Eudract [0.31,0.32] [0.31,0.32] 0.36 0.37 0.32
Flight [0.66,0.71] [0.54,0.54] 0.68 0.79 0.58
$F$ Hospital [0.88,0.90] [0.16,0.17] 0.92 0.92 n/a
Allergen [0.00,0.55] [0.41,0.41] 0.13 0.13 0.41
Eudract [0.44,0.47] [0.44,0.47] 0.51 0.52 0.27
Flight [0.67,0.72] [0.65,0.65] 0.69 0.79 0.57
Table 4 provides an overview of the repair quality obtained by the two configurations of Llunatic and the three configurations of Swipe on all data sets. For each combination of data set and measure of quality, we indicate the best performing approach in bold font. For recall and $F$-score on the Allergen data set, we did not indicate a best approach as it cannot be indicated with certainty. When inspecting the results in Table 4, several interesting observations can be made. A first observation is that results are sensitive to the exact configurations. This is especially true for the Allergen dataset, which provides allergen information for each product coming from only two sources. This low number of sources makes voting procedures less suitable. The resolution of conflicts by means of the $\mathsf{max}$ repair function, is a superior choice here. This behaviour is also encoded in the FT scenario of Llunatic and it was confirmed that in this case, Llunatic and Swipe produced exactly the same results.
Interestingly, for other datasets, the voting scenarios of Swipe provide robust and good results. When a data set has no null values (Hospital and Allergen), both voting scenarios are equivalent. When null values are present (Eudract and Flight), weighted voting produces better results and turns out to be a good baseline choice. The $\mathsf{max}$ function provide good results only in particular scenarios like the Allergen dataset.
For Llunatic, the sensitivity to configuration choices seems to be higher than for Swipe. For the Hospital data set, there is a large difference in repair quality between the two configurations, most likely due to the usage of backward repairing in the second configuration. In general, we found that giving Llunatic the possibility to include backward repairs (this is the case in the fine-tuned configurations) seems to decrease repair quality on the data sets considered here. Generalizing this observation should be done with caution, but the results reported here make a strong case for at least questioning the usefulness of backward repair on real-life data sets.
@@ 5.4 Scalability
In this section, we investigate the run time efficiency and scalability of Swipe. First, we measure the mean run time over five runs of Swipe and Llunatic on the different data sets and report the mean execution times (in seconds). We do not report other scalability parameters like CPU usage or main memory consumption, although we have investigated these parameters and came to similar conclusions. In case of Llunatic, we differentiate between the standard configuration and the configuration that was fine-tuned for each data set (Table 3). For Swipe, we took for each dataset, the best performing repair function. The results are reported in Table 5.
@@
Data set Llunatic (S) Llunatic (FT) Swipe (best)
@@
The results from Table 5 show that Swipe outperforms Llunatic on each data set in terms of execution time. This comes as no surprise as Swipe is designed explicitly as a simplification of the multi-sequence Chase algorithm that underlies Llunatic. Yet, we observe that the gain in execution time is considerable. Swipe is two to three orders of magnitude faster than the fine-tuned configuration of Llunatic. When we compare to the standard Llunatic configuration, the difference reduces to one order of magnitude. This provides us with some interesting insights. First of all, the cost manager of Llunatic does what it is supposed to do: restricting repair operations to forward repair only and limiting the branching factor significantly reduces the execution time of Llunatic. However, Table 4 shows that this gain in computational efficiency also leads to lesser quality of the repairs. In that sense, the Swipe algorithm makes more considerate choices in reducing computational cost with better repair quality as a result.